When (8x + 6) (ax + b) = (32x^2) - 40x -48. Find a and b
Answer by ankor@dixie-net.com(22740)

When (8x + 6) (ax + b) = (32x^2) - 40x - 48. Find a and b
FOIL (8x+6)(ax+b)
8ax^2 + 8bx + 6ax + 6b
then
8ax^2 = 32x^2
Divide both sides by 8x^2
a = 4
:
Find b
6b = -48
b = -48/6
b = -8
:
Check, FOIL
(8x+6)*(4x-8) = 32x^2 - 64x + 24x - 48 = 32x^2 - 40x - 48